# Self-consistency error correction for accurate machine learning potentials from variational Monte Carlo

### v1

@@ A.2 Cutoff in the correction error
[…]
To manage these erratic components, in this study, we retained the biased force values in the corrected data set whenever three standard deviations exceeded a given threshold, set here to $0.015$ Ha/Bohr. As shown in Fig. 6, this threshold corresponds to a minimum in the standard deviation distribution for the corrected forces, indicating the onset of departure from normality. A similar strategy was employed for pressure, using a threshold of $1.5\times 10^{-5}$ a.u. We confirmed that forces and pressures with error bars below the threshold are unbiased, consistent with their corresponding potential energy surfaces (PESs), as illustrated in Figs.1 and 2. This choice was further validated by analyzing the behavior of $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ (see Eq.(12)) as a function of the relative weight $W_{F}/W_{E}$ in the loss function, using models trained on datasets corrected with different thresholds for $\sigma_{F}$. The results, depicted in Fig. 7, confirm that the threshold of $0.015$ Ha/Bohr yields the best overall performance. Indeed, although $\Delta_{E}$ and $\Delta_{P}$ both steadily decrease with increasing threshold values (for fixed $W_{F}/W_{E}$), the accuracy of forces deteriorates significantly when corrections with large error bars (corresponding to the largest threshold, i.e. $3\sigma_{F}\leq 0.04$ Ha/Bohr) are included.
@@ A.3 MACE models comparison
Here, we further support the conclusions made in Sec. 3.2 by carrying out the same analysis using MACE 35. This is a completely different MLIP architecture, because it is a framework based on equivariant message-passing neural networks with high-body order messages. The scope of this complementary analysis is to demonstrate that the behavior observed with the KRR models is general and only due to the training set consistency, and not linked to any particular implementation of the MLIP. The models consider $256$ inner invariant features and a cutoff of $3$ (effectively doubled when considering the message passing step). For each dataset, both training and test sets are the same as the one considered before for the KRR models in Sec. 3.2. The Adam optimizer 60 is used to find the models parameters. In particular, at each step the gradient of the loss function is evaluated on a subset of the training set ("batch") of $N_{b}=4$ configurations. The results for the different datasets are reported in Fig. 8, where we show the behavior of $\Delta_{X}$, defined in Eq. (12), for energy, forces, and pressure.
@@ A.4 Molecular dynamics simulation details
At each step, the energy, forces, and pressure were calculated at the DFT level using the Quantum Espresso package in its GPU accelerated version 61, 62, 63 with the PBE functional, and then summed with those predicted by the different MLIPs. For the DFT simulations, a $60$ Ry plane-wave cutoff with a projector augmented wave pseudopotential 64 was used together with a $4\times 4\times 4$ Monkhorst-Pack $\mathbf{k}$-point grid. For the MD simulations we used a time step of $0.25$ fs and a Langevin thermostat 65, 41 with damping $\gamma=0.13$ fs-1.

### v2

@@ A.2 Cutoff in the correction error
Figure 6 shows the error bars of the corrected forces in the dataset. We observed that, for certain force components and pressures ($\sim 10\%$ of all components), the correction is occasionally accompanied by large error bars. We identified the cause of this phenomenon in our current workflow, which estimates the parameter derivatives $\cfrac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}i}$ and $\cfrac{d\lambda_{k,k^{\prime}}}{dV}$ in Eq.10 and Eq.11 using the FDM by performing multiple DFT calculations at displaced coordinates Nakano2024. On the one hand, when employing localized basis sets in DFT, a cutoff is usually applied to reduce basis redundancy by eliminating elements corresponding to overlap matrix eigenvalues below a certain threshold, as written in Sec. 2.2. Without this cutoff, the forces and pressures computed according to Eq. 5 exhibit large error bars, as reported in Ref.Nakano2021. On the other hand, applying this cutoff causes the basis set to depend on atomic displacements in the FDM approach, as the quality of the basis set changes when ions move. This displacement-induced dependence occasionally deteriorates the error cancellation in Eqs. 10 and 11, resulting in larger error bars. Another source of large error bars is the finite smearing parameter used for molecular orbital (MO) occupations. Since hydrogen exhibits metallic behavior depending on its density, a smearing technique is necessary to achieve stable and smooth convergence for all configurations. However, employing smearing methods can introduce discontinuities when evaluating parameter derivatives $\cfrac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}i}$ and $\cfrac{d\lambda_{k,k^{\prime}}}{dV}$ by FDM. Additionally, the discrete real-space grid size used in DFT calculations contributes to error deterioration. Within the FDM approach, we need to perform 3NN DFT calculations to compute the correction terms. To keep their computational cost low for large $N$, we chose a real-space grid size of $0.15^{3}$ Bohr3. We underline that the finite-difference step in ionic positions and cell volume for FDM, the smearing parameter, and the real-space grid have been chosen to provide converged forces and pressures, such that any possible bias is comparable with their total stochastic error bars. Yet, the occurrence of large error bars in the SCE correction could not be avoided for some particular ionic configurations by playing solely with these parameters. These factors indicate that the FDM method proposed in Ref. Nakano2024 can be improved for metallic systems in the future, although it is already robust for insulating systems with band gaps, as we have verified for c-BN, SiC, Si, and diamond.
To manage these erratic components, in this study, we retained the biased force values in the corrected data set whenever three standard deviations exceeded a given threshold, set here to $0.015$ Ha/Bohr. As shown in Fig. 6, this threshold corresponds to a minimum in the standard deviation distribution for the corrected forces, indicating the onset of departure from normality. A similar strategy was employed for pressure, using a threshold of $1.5\times 10^{-5}$ a.u. We confirmed that forces and pressures with error bars below the threshold are unbiased, consistent with their corresponding potential energy surfaces (PESs), as illustrated in Figs.1 and 2. This choice was further validated by analyzing the behavior of $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ (see Eq.(12)) as a function of the relative weight $W_{F}/W_{E}$ in the loss function, using models trained on datasets corrected with different thresholds for $\sigma_{F}$. The results, depicted in Fig. 7, confirm that the threshold of $0.015$ Ha/Bohr yields the best overall performance. Indeed, although $\Delta_{E}$ and $\Delta_{P}$ both steadily increase with increasing threshold values (for fixed $W_{F}/W_{E}$), the accuracy of forces deteriorates significantly when corrections with large error bars (corresponding to the largest threshold, i.e. $3\sigma_{F}\leq 0.04$ Ha/Bohr) are included.
@@ A.3 MACE models comparison
Here, we further support the conclusions made in Sec. 3.2 by carrying out the same analysis using MACE Batatia2022. This is a completely different MLIP architecture, because it is a framework based on equivariant message-passing neural networks with high-body order messages. The scope of this complementary analysis is to demonstrate that the behavior observed with the KRR models is general and only due to the training set consistency, and not linked to any particular implementation of the MLIP. The models consider $256$ inner invariant features and a cutoff of $3$ Å (effectively doubled when considering the message passing step). For each dataset, both training and test sets are the same as the one considered before for the KRR models in Sec. 3.2. The Adam optimizer Kingma2014 is used to find the models parameters. In particular, at each step the gradient of the loss function is evaluated on a subset of the training set ("batch") of $N_{b}=4$ configurations. The results for the different datasets are reported in Fig. 8, where we show the behavior of $\Delta_{X}$, defined in Eq. (12), for energy, forces, and pressure.
@@ A.4 Molecular dynamics simulation details
At each step, the energy, forces, and pressure were calculated at the DFT level using the Quantum Espresso package in its GPU accelerated version Giannozzi2009, Giannozzi2017, Giannozzi2020 with the PBE functional, and then summed with those predicted by the different MLIPs. For the DFT simulations, a $60$ Ry plane-wave cutoff with a projector augmented wave pseudopotential PAWpseudo was used together with a $4\times 4\times 4$ Monkhorst-Pack $\mathbf{k}$-point grid. For the MD simulations we used a time step of $0.25$ fs and a Langevin thermostat Ricci2003, Attaccalite2008 with damping $\gamma=0.13$ fs-1.